### v3

@@ Abstract
Modern medical records include a vast amount of multimodal free text clinical data and imaging data from radiology, cardiology, and digital pathology. Fully mining such big data requires multitasking; otherwise, occult but important aspects may be overlooked, adversely affecting clinical management and population healthcare. Despite remarkable successes of AI in individual tasks with single-modal data, the progress in developing generalist medical AI (GMAI) remains relatively slow to combine multimodal data for multitasks because of the dual challenges of data curation and model architecture. The data challenge involves querying and curating multimodal structured and unstructured text, alphanumeric, and especially 3D tomographic scans on an individual patient level for real-time decisions and on a scale to estimate population health statistics. The model challenge demands a scalable and adaptable network architecture to integrate multimodal datasets for diverse clinical tasks. Here we propose the first-of-its-kind medical multimodal-multitask foundation model (M3FM) with the initial application in low-dose CT lung cancer screening and related tasks. After we curated a comprehensive multimodal multitask dataset consisting of 49 clinical data types including 163,725 chest CT series and 17 medical tasks involved in lung cancer screening, we develop a multimodal question-answering framework as a unified training and inference strategy to synergize multimodal information and perform multiple tasks via free-text prompting. Our M3FM consistently and significantly outperforms the state-of-the-art single-modal task-specific models, improving lung cancer risk prediction by up to 20% and cardiovascular disease mortality risk estimation by up to 10%. M3FM identifies multimodal data elements informative for clinical tasks, being instrumental in gaining insights and correlating the multimodal data with relevant diseases. M3FM can flexibly adapt to new tasks with a small out-of-distribution dataset, effectively handle various combinations of multimodal data, and efficiently process high-dimensional images at multiple scales. In a broader sense, as a specialty-oriented generalist medical AI (SOGMAI) model, M3FM innovates lung cancer management and related tasks. This SOGMAI approach paves the way for similar breakthroughs in other areas of medicine, closing the gap between specialists and the generalist.
@@ 1 Introduction
In this study, we report the first-of-its-kind SOGMAI model, M3FM, that can perceive multimodal data to multiple tasks dedicated to LCS, as illustrated in Figure 1. The success of our M3FM for LCS is achieved by addressing both the data and model challenges.
First, we present an integrated and scalable data curation approach to align high-dimensional medical images with other clinical datasets for specialty-oriented tasks by leveraging domain-specific expertise. According to the LCS practice, 17 tasks of interest were defined and 49 different data elements were involved. To collect the multimodal data for these tasks, we were granted access to massive data from the National Lung Cancer Screening Trial (NLST) and the Medical Imaging and Data Resource Center (MIDRC), both of which involved multiple medical institutes. Also, we independently collected two similar datasets from Wake Forest University School of Medicine (WFUSM) and Massachusetts General Hospital (MGH). Our datasets were curated by aligning among a variety of data types for each task, including chest CT scans and the associated imaging parameters, radiology reports, de-identified demographics, smoking history, disease history, cancer history, family history, pathological results, and follow-up data. The resulting 64 sub-datasets for training, validating, testing, and fine-tuning are summarized in Figure 2 and Tabel 1. It is worth noting that we used only medical records without expensive human labeling, such as bounding box annotations for lung nodules or other imaging findings, during the data curation so that these datasets could be easily enlarged.
@@ 2.1 Multimodal Multitask Datasets
Figure 2-(a) shows the general data curation pipeline, including medical tasks of interest definition, task-specific multimodal data collection, multimodal data processing and alignment, and MQA dataset construction. We target 17 (sub-)tasks related to the LCS process, including 5 tasks for lung nodule detection and characterization, 1 task for cardiovascular disease (CVD) diagnosis, 1 task for CVD mortality prediction, 1 task for lung cancer risk prediction over multiple years, 7 tasks for other chest abnormality exams, 1 task for COVID-19 detection, as well as 1 task for American College of Radiology (ACR) guidelines for Lung CT Screening Reporting and Data System (Lung-RADS) categorization. COVID-19 detection from CT is included since it remains a global threat [42]. The ground-truth labels come from different information sources, including radiology reports, disease history, pathology test results, follow-up data, death reports, and laboratory test results for different tasks as described in Table 1. To curate the multimodal datasets, multiple data sources are aligned, including volumetric CT scans, demographics, smoking history, disease history, cancer history, family cancer history, and other task-specific clinical data. In total, 49 different clinical data types are involved in the multimodal datasets for LCS, as described in Table 1. For each task, one training, one validation, and one or more testing datasets were constructed. An overview of our multimodal multitask dataset is given in Figure 2-(b). The data were collected from different data centers and institutes, including NLST, MIDRC, Wake Forest University School of Medicine (WFUSM), and Massachusetts General Hospital (MGH). In total, we curated 17 training, 17 validation, and 29 testing datasets for the 17 tasks, with the detailed composition of each dataset in Figure 2-(c), (d), and (e). We also collected an out-of-distribution multimodal dataset from WFUSM for transfer learning. To inspect the modeling ability for textural clinical data, we simulated a dataset for clinical information retrieval, as illustrated in Figure 1. Since we unify multitask learning with multimodal data in an MQA framework, each dataset consists of task-specific multimodal inputs, questions, and answers. The details for all tasks are summarised in Table 1.
@@
As the first data source, we were granted access to all recorded data in NLST NLST is a randomized trial for evaluating LCS with 3D LDCT versus 2D chest radiography, demonstrating that screening with LDCT lowered lung cancer mortality by 20%. The NLST data were collected from 33 medical institutions, which were randomly indexed without revealing their identifications publicly. The 26,722 participants in the LDCT screening arm were enrolled from August 2002 through April 2004. The participants underwent three screenings at 1-year intervals from August 2002 through September 2007. The follow-up data were collected until December 31, 2009. During the whole process, diverse data were recorded, including demographics, smoking history, disease history, multiple CT series with different reconstruction algorithms and associated imaging parameters, key abnormalities in structured reports, pathology test results for lung cancer, follow-up data, and vital status. Being consistent with clinical practice, we constructed 15 multimodal datasets for 15 tasks, including 5 datasets for predicting the presence of lung nodules and estimating the location, size, margin, and attenuation properties of lung nodules; 7 datasets for identifying chest abnormalities, including atelectasis, pleural thickening/effusion, non-calcified hilar/mediastinal adenopathy/mass, chest wall abnormality (bone destruction, metastasis, etc.), consolidation, emphysema, reticular/reticulonodular opacities/honeycombing/fibrosis/scar; 1 dataset for cardiovascular abnormality diagnosis; 1 dataset for cardiovascular mortality risk prediction; and 1 dataset for lung cancer risk prediction within from 1 to 6 years. Each dataset was randomly split into training, validation, and test datasets. The patient-wise information in the validation and test datasets is not leaked to the training datasets across all tasks. From NLST, we included 125,090 effective volumetric chest CT scans of 26,254 patients.
@@ 2.2 M3FM Setting New Records
Figure 3-(a) summarizes the key results of M3FM against the SOTA models [33, 35] on the OpenM3Chest dataset. Overall, M3FM sets new records for all tasks, significantly outperforming the SOTA models for most of these tasks. For a fair comparison, we retrained the Sybil model [33], denoted as Sybil∗, for lung cancer risk prediction without using the costly bounding box annotations but predicting lung cancer risks by merging the separate results of left and right lungs. It is observed that Sybil∗ achieved inferior results for 12-year risk prediction but superior results for 36-year risk prediction in comparison with the results obtained using the original Sybil model. Without using any bounding box, our M3FM achieves an Area Under the receiver operating characteristic Curve (AUC) of 94.0 (95% 95 CI, 91.19-96.98), 88.81 (95% CI, 85.67-91.95), 85.99 (95% CI, 82.88-89.1), 86.04 (95% CI, 83.1-88.98), 83.92 (95% CI, 80.98-86.85), 82.32 (95% CI, 79.36-85.29) for lung cancer risk prediction over six years, outperforming both Sybil∗ and original Sybil models by the margins of 5% to 9% and 2% to 11%, respectively. For CVD diagnosis and CVD mortality prediction, we compared the results on both the original dataset [35] and our OpenM3Chest dataset. M3FM achieves an AUC of 92.84 (95% CI, 91.36-94.33) for CVD diagnosis and an AUC of 89.04 (95% CI, 84.27-93.81) for CVD mortality prediction on the OpenChest dataset, outperforming the previous models (Tri2D-Net [35]) by 5% and 9% respectively, and achieves an AUC of 92.03% for CVD diagnosis and 86.14% for CVD mortality prediction on the datasets constructed in [35], outperforming the previous models (Tri2D-Net) by 5% and 10% respectively. For several tasks including nodule detection, nodule localization, nodule size prediction, and emphysema detection, M3FM improves the results by various degrees up to 3% of AUC. For all the other tasks, M3FM significantly improves the performance from 5% to 10%. To study the scalability of M3FM, we trained three versions of M3FM, consisting of 257M (M3FM-Base), 502M (M3FM-Large), and 865M (M3FM-Huge) trainable parameters respectively. The results of these three models are summarized in Figure 3-(b). Overall, with a larger model size, the performance becomes better, especially from M3FM-Base to M3FM-Large. This trend is consistent with the well-known scaling law in the field of foundation models.
@@ 2.3 M3FM Encoding Multimodal Data and Synergizing Multiple Clinical Tasks
Table 2 compares the results of the single-modality single-task, multi-modality single-task, and multi-modality multitask M3FM-Large models. First, the single-modality single-task models were trained and evaluated on LDCT data only and denoted by M3FM-SM-ST, while the multi-modality and single-task models were trained and evaluated on multimodal data and denoted by M3FM-MM-ST. Overall, the multimodal information improves the prediction results for multiple tasks. In particular, M3FM-ST-SM-ST achieves an AUC of 81.63 (95% CI, 75.85-87.41) for CVD mortality prediction while the M3FM-MM-ST model achieves an AUC of 87.09 (95% CI, 82.0-92.19), which represents a 5.46% improvement. While M3FM-SM-ST achieves an AUC of 89.24 (95% CI, 87.45-91.04) for CVD diagnosis, the M3FM-MM-ST model achieves an AUC of 92.38 (95% CI, 90.84-93.92), i.e., a 3.14% improvement. Similarly, M3FM-SM-ST achieves an AUC of 65.15 (95% CI, 59.39-70.92) for consolidation detection, and the M3FM-MM-ST model achieves an AUC of 68.95 (95% CI, 63.26-74.64), a 3.80% improvement. Also, M3FM-SM-ST achieves an AUC of 76.76 (95% CI, 75.73-77.79) for reticular/reticulonodular opacities/honeycombing/fibrosis/scar detection, and the M3FM-MM-ST model achieves an AUC of 79.29 (95% CI, 78.3-80.27), a 2.53% improvement. It is further observed that M3FM-MM-ST models produce slightly improved or comparable results in comparison with M3FM-SM-ST for the other tasks. Then, we compared the multimodal multitask model (M3FM-MM-MT) and multimodal single-task models (M3FM-MM-ST). Impressively, training on multiple tasks, M3FM-MT-MM outperforms the M3FM-ST-MM for 17 out of 22 (sub)-tasks.
@@ 2.4 M3FM Identifying Clinically Informational Elements
Then, we evaluated if M3FMs could effectively encode the physical size information. The ablation results in Table 5 show that the embedded physical size information of LDCT improves the AUC results for multiple tasks. The physical size information boosts the AUC of 16-year lung cancer risk prediction by 5%, 4%, 4%, 7%, 8% and 12%, respectively. The physical size information also improves AUC results of the nodule size characterization, CVD diagnosis, and CVD mortality prediction, by 0.71%, 0.47%, and 1.11% respectively.
@@ 2.5 M3FM Improving Generalizability
We evaluated the generalizability of M3FMs on the independently collected datasets, with the comparison results on the MGH and WFUSM datasets shown in Table 6. For the CVD diagnosis task, we constructed two datasets, which regard moderate and severe CVD as positive and severe CVD only as positive, respectively. On the two MGH CVD datasets, the multimodal multitask model (M3FM-MM-MT) improved the AUC by 10.60% and 6.57% relative to the previous model, improved the AUC by 4.85% and 2.36% relative to the single-modality single-task model (M3FM-SM-ST), and also achieved slight AUC improvements relative to the multi-modality single-task model (M3FM-MM-ST). Relative to M3FM-SM-ST, the M3FM-MM-ST model improved the AUC by 4.39% and 1.75% on the two CVD MGH datasets. For the 1-year lung cancer risk prediction on the MGH dataset, the M3FM-MM-MT model improved the AUC by 20.80% than the previous model under the same experimental settings without using any bounding box annotations, improved the AUC by 4.85% than M3FM-SM-ST, and improved the AUC by 6.89% than M3FM-MM-ST. On the MGH emphysema, atelectasis, and reticular opacities/honeycombing/fibrosis/scar datasets, the M3FM improved the AUC by 5.23%, 14.34%, and 12.91% than the previous model, and also achieved AUC improvements by 0.24% 4.96% than M3FM-SM-ST and M3FM-MM-ST. For the CVD tasks on MGH datasets, M3FM-MM-MT improved the AUC by 12% and 6.29% than the previous model, improved the AUC by 6.46% and 3.77% relative to M3FM-SM-ST, and had the same results with M3FM-MM-ST. For the 1-year lung cancer risk prediction on the MGH dataset, the M3FM-MM-MT model improved the AUC by 18.54% relative to the previous model under the same experimental settings without using any bounding box annotations, improved the AUC by 5.91% than M3FM-SM-ST, and improved the AUC by 2.57% than M3FM-MM-ST; and M3FM-MM-ST improved the AUC by 3.24% than M3FM-SM-ST. On the WFUSM emphysema, atelectasis, and reticular opacities/honeycombing/fibrosis/scar datasets, the M3FM-MM-MT model improved the AUC by 0.78%, 9.65%, and 14.24% than the previous model.

### v4

@@ Abstract
Modern medical records include a vast amount of multimodal free text clinical data and imaging data from radiology, cardiology, and digital pathology. Fully mining such big data requires multitasking; otherwise, occult but important aspects may be overlooked, adversely affecting clinical management and population healthcare. Despite remarkable successes of AI in individual tasks with single-modal data, the progress in developing generalist medical AI (GMAI) remains relatively slow to combine multimodal data for multitasks because of the dual challenges of data curation and model architecture. The data challenge involves querying and curating multimodal structured and unstructured text, alphanumeric, and especially 3D tomographic scans on an individual patient level for real-time decisions and on a scale to estimate population health statistics. The model challenge demands a scalable and adaptable network architecture to integrate multimodal datasets for diverse clinical tasks. Here we propose the first-of-its-kind medical multimodal-multitask foundation model (M3FM) with application in low-dose CT lung cancer screening and related tasks. After we curated a comprehensive multimodal multitask dataset consisting of 49 clinical data types including 163,725 chest CT series and 17 medical tasks involved in lung cancer screening, we develop a multimodal question-answering framework as a unified training and inference strategy to synergize multimodal information and perform multiple tasks via free-text prompting. Our M3FM consistently and significantly outperforms the state-of-the-art single-modal task-specific models, improving lung cancer risk prediction by up to 20% and cardiovascular disease mortality risk estimation by up to 10%. M3FM identifies multimodal data elements informative for clinical tasks, being instrumental in gaining insights and correlating the multimodal data with relevant diseases. M3FM can flexibly adapt to new tasks with a small out-of-distribution dataset, effectively handle various combinations of multimodal data, and efficiently process high-dimensional images at multiple scales. In a broader sense, as a specialty-oriented generalist medical AI (SOGMAI) model, M3FM innovates lung cancer management and related tasks. This SOGMAI approach paves the way for similar breakthroughs in other areas of medicine, closing the gap between specialists and the generalist.
@@ 1 Introduction
In this study, we report the first-of-its-kind SOGMAI model, M3FM, that can perceive multimodal data to perform multiple tasks dedicated to LCS, as illustrated in Figure 1. The success of our M3FM for LCS is achieved by addressing both the data and model challenges.
First, we present an integrated and scalable data curation approach to align high-dimensional medical images with other clinical datasets for specialty-oriented tasks by leveraging domain-specific expertise. According to the LCS practice, 17 tasks of interest were defined and 49 different data elements were involved. To collect the multimodal data for these tasks, we were granted access to massive data from the National Lung Cancer Screening Trial (NLST) [22] and the Medical Imaging and Data Resource Center (MIDRC) [39], both of which involved multiple medical institutes. Also, we independently collected two similar datasets from Wake Forest University School of Medicine (WFUSM) and Massachusetts General Hospital (MGH). Our datasets were curated by aligning among a variety of data types for each task, including chest CT scans and the associated imaging parameters, radiology reports, de-identified demographics, smoking history, disease history, cancer history, family history, pathological results, and follow-up data. The resulting 64 sub-datasets for training, validating, testing, and fine-tuning are summarized in Figure 2 and Tabel 1. It is worth noting that we used only medical records without expensive human labeling, such as bounding box annotations for lung nodules or other imaging findings, during the data curation so that these datasets could be easily enlarged.
@@ 2.1 Multimodal Multitask Datasets
Figure 2 (a) shows the general data curation pipeline, including medical tasks of interest definition, task-specific multimodal data collection, multimodal data processing and alignment, and MQA dataset construction. We target 17 (sub-)tasks related to the LCS process, including 5 tasks for lung nodule detection and characterization, 1 task for cardiovascular disease (CVD) diagnosis, 1 task for CVD mortality prediction, 1 task for lung cancer risk prediction over multiple years, 7 tasks for other chest abnormality exams, 1 task for COVID-19 detection, as well as 1 task for American College of Radiology (ACR) guidelines for Lung CT Screening Reporting and Data System (Lung-RADS) categorization. COVID-19 detection from CT is included since it remains a global threat [43]. The ground-truth labels come from different information sources, including radiology reports, disease history, pathology test results, follow-up data, death reports, and laboratory test results for different tasks as described in Table 1. To curate the multimodal datasets, multiple data sources are aligned, including volumetric CT scans, demographics, smoking history, disease history, cancer history, family cancer history, and other task-specific clinical data. In total, 49 different clinical data types are involved in the multimodal datasets for LCS, as described in Table 1. For each task, one training, one validation, and one or more testing datasets were constructed. An overview of our multimodal multitask dataset is given in Figure 2 (b). The data were collected from different data centers and institutes, including NLST, MIDRC, Wake Forest University School of Medicine (WFUSM), and Massachusetts General Hospital (MGH). In total, we curated 17 training, 17 validation, and 29 testing datasets for the 17 tasks, with the detailed composition of each dataset in Figure 2 (c), (d), and (e). We also collected an out-of-distribution multimodal dataset from WFUSM for transfer learning. To inspect the modeling ability for textural clinical data, we simulated a dataset for clinical information retrieval, as illustrated in Figure 1. Since we unify multitask learning with multimodal data in an MQA framework, each dataset consists of task-specific multimodal inputs, questions, and answers. The details for all tasks are summarised in Table 1.
@@
As the first data source, we were granted access to all recorded data in NLST. NLST is a randomized trial for evaluating LCS with 3D LDCT versus 2D chest radiography, demonstrating that screening with LDCT lowered lung cancer mortality by 20%. The NLST data were collected from 33 medical institutions, which were randomly indexed without revealing their identifications publicly. The 26,722 participants in the LDCT screening arm were enrolled from August 2002 through April 2004. The participants underwent three screenings at 1-year intervals from August 2002 through September 2007. The follow-up data were collected until December 31, 2009. During the whole process, diverse data were recorded, including demographics, smoking history, disease history, multiple CT series with different reconstruction algorithms and associated imaging parameters, key abnormalities in structured reports, pathology test results for lung cancer, follow-up data, and vital status. Being consistent with clinical practice, we constructed 15 multimodal datasets for 15 tasks, including 5 datasets for predicting the presence of lung nodules and estimating the location, size, margin, and attenuation properties of lung nodules; 7 datasets for identifying chest abnormalities, including atelectasis, pleural thickening/effusion, non-calcified hilar/mediastinal adenopathy/mass, chest wall abnormality (bone destruction, metastasis, etc.), consolidation, emphysema, reticular/reticulonodular opacities/honeycombing/fibrosis/scar; 1 dataset for cardiovascular abnormality diagnosis; 1 dataset for cardiovascular mortality risk prediction; and 1 dataset for lung cancer risk prediction within from 1 to 6 years. Each dataset was randomly split into training, validation, and test datasets. The patient-wise information in the validation and test datasets is not leaked to the training datasets across all tasks. From NLST, we included 125,090 effective volumetric chest CT scans of 26,254 patients.
@@ 2.2 M3FM Setting New Records
Figure 3 (a) summarizes the key results of M3FM against the SOTA models [33, 35] on the OpenM3Chest dataset. Overall, M3FM sets new records for all tasks, significantly outperforming the SOTA models for most of these tasks. We used the Area Under the receiver operating characteristic Curve (AUC) and the 95% two-sided Confidence Intervals (CI) of AUC values proposed by Hanley and McNeil [44] as the evaluation metrics. For a fair comparison, we retrained the Sybil model [33], denoted as Sybil∗, for lung cancer risk prediction without using the costly bounding box annotations but predicting lung cancer risks by merging the separate results of left and right lungs. It is observed that Sybil∗ achieved inferior results for 12-year risk prediction but superior results for 36-year risk prediction in comparison with the results obtained using the original Sybil model. Without using any bounding box, our M3FM achieves an AUC of 94.00% (95% CI, 91.19-96.98), 88.81% (95% CI, 85.67%-91.95%), 85.99% (95% CI, 82.88%-89.1%), 86.04% (95% CI, 83.10%-88.98%), 83.92% (95% CI, 80.98%-86.85%), 82.32% (95% CI, 79.36%-85.29%) for lung cancer risk prediction over six years, outperforming both Sybil∗ and original Sybil models by the margins of 5% to 9% and 2% to 11%, respectively. For CVD diagnosis and CVD mortality prediction, we compared the results on both the original dataset [35] and our OpenM3Chest dataset. M3FM achieves an AUC of 92.84% (95% CI, 91.36%-94.33%) for CVD diagnosis and an AUC of 89.04% (95% CI, 84.27%-93.81%) for CVD mortality prediction on the OpenChest dataset, outperforming the previous models (Tri2D-Net [35]) by 5% and 9% respectively, and achieves an AUC of 92.03% for CVD diagnosis and 86.14% for CVD mortality prediction on the datasets constructed in [35], outperforming the previous models (Tri2D-Net) by 5% and 10% respectively. For several tasks including nodule detection, nodule localization, nodule size prediction, and emphysema detection, M3FM improves the results by various degrees up to 3% of AUC. For all the other tasks, M3FM significantly improves the performance from 5% to 10%. To study the scalability of M3FM, we trained three versions of M3FM, consisting of 257M (M3FM-Base), 502M (M3FM-Large), and 865M (M3FM-Huge) trainable parameters respectively. The results of these three models are summarized in Figure 3 (b). Overall, with a larger model size, the performance becomes better, especially from M3FM-Base to M3FM-Large. This trend is consistent with the well-known scaling law [20] in the field of foundation models.
@@ 2.3 M3FM Encoding Multimodal Data and Synergizing Multiple Clinical Tasks
Table 2 compares the results of the single-modality single-task, multi-modality single-task, and multi-modality multitask M3FM-Large models. First, the single-modality single-task models were trained and evaluated on LDCT data only and denoted by M3FM-SM-ST, while the multi-modality and single-task models were trained and evaluated on multimodal data and denoted by M3FM-MM-ST. Overall, the multimodal information improves the prediction results for multiple tasks. In particular, M3FM-SM-ST achieves an AUC of 81.63% (95% CI, 75.85%-87.41%) for CVD mortality prediction while the M3FM-MM-ST model achieves an AUC of 87.09% (95% CI, 82.00%-92.19%), which represents a 5.46% improvement. While M3FM-SM-ST achieves an AUC of 89.24% (95% CI, 87.45%-91.04%) for CVD diagnosis, the M3FM-MM-ST model achieves an AUC of 92.38% (95% CI, 90.84%-93.92%), i.e., a 3.14% improvement. Similarly, M3FM-SM-ST achieves an AUC of 65.15% (95% CI, 59.39%-70.92%) for consolidation detection, and the M3FM-MM-ST model achieves an AUC of 68.95% (95% CI, 63.26%-74.64%), a 3.80% improvement. Also, M3FM-SM-ST achieves an AUC of 76.76% (95% CI, 75.73%-77.79%) for reticular/reticulonodular opacities/honeycombing/fibrosis/scar detection, and the M3FM-MM-ST model achieves an AUC of 79.29% (95% CI, 78.30%-80.27%), a 2.53% improvement. It is further observed that M3FM-MM-ST models produce slightly improved or comparable results in comparison with M3FM-SM-ST for the other tasks. Then, we compared the multimodal multitask model (M3FM-MM-MT) and multimodal single-task models (M3FM-MM-ST). Impressively, training on multiple tasks, M3FM-MT-MM outperforms the M3FM-ST-MM for 17 out of 22 (sub)-tasks.
@@ 2.4 M3FM Identifying Clinically Informational Elements
Then, we evaluated if M3FMs could effectively encode the physical size information. The ablation results in Figure 4 show that the embedded physical size information of LDCT improves the AUC results for multiple tasks. The physical size information boosts the AUC of 16-year lung cancer risk prediction by 5%, 4%, 4%, 7%, 8% and 12%, respectively. The physical size information also improves AUC results of the nodule size characterization, CVD diagnosis, and CVD mortality prediction, by 0.71%, 0.47%, and 1.11% respectively.
@@ 2.5 M3FM Improving Generalizability
We evaluated the generalizability of M3FMs on the independently collected datasets, with the comparison results on the MGH and WFUSM datasets shown in Figure 6. For the CVD diagnosis task, we constructed two datasets, which regard moderate and severe CVD as positive and severe CVD only as positive, respectively. On the two MGH CVD datasets, the multimodal multitask model (M3FM-MM-MT) improved the AUC by 10.60% and 6.57% relative to the previous model, improved the AUC by 4.85% and 2.36% relative to the single-modality single-task model (M3FM-SM-ST), and also achieved slight AUC improvements relative to the multi-modality single-task model (M3FM-MM-ST). Relative to M3FM-SM-ST, the M3FM-MM-ST model improved the AUC by 4.39% and 1.75% on the two CVD MGH datasets. For the 1-year lung cancer risk prediction on the MGH dataset, the M3FM-MM-MT model improved the AUC by 20.80% than the previous model under the same experimental settings without using any bounding box annotations, improved the AUC by 4.85% than M3FM-SM-ST, and improved the AUC by 6.89% than M3FM-MM-ST. On the MGH emphysema, atelectasis, and reticular opacities/honeycombing/fibrosis/scar datasets, the M3FM improved the AUC by 5.23%, 14.34%, and 12.91% than the previous model, and also achieved AUC improvements by 0.24% 4.96% than M3FM-SM-ST and M3FM-MM-ST. For the CVD tasks on MGH datasets, M3FM-MM-MT improved the AUC by 12% and 6.29% than the previous model, improved the AUC by 6.46% and 3.77% relative to M3FM-SM-ST, and had the same results with M3FM-MM-ST. For the 1-year lung cancer risk prediction on the MGH dataset, the M3FM-MM-MT model improved the AUC by 18.54% relative to the previous model under the same experimental settings without using any bounding box annotations, improved the AUC by 5.91% than M3FM-SM-ST, and improved the AUC by 2.57% than M3FM-MM-ST; and M3FM-MM-ST improved the AUC by 3.24% than M3FM-SM-ST. On the WFUSM emphysema, atelectasis, and reticular opacities/honeycombing/fibrosis/scar datasets, the M3FM-MM-MT model improved the AUC by 0.78%, 9.65%, and 14.24% than the previous model.